Solve (x-1)^5 = 12
Answer by ankor@dixie-net.com(22740)

Solve (x-1)^5 = 12
:
Take the fifth root of both sides and you have:
x - 1 = 12^(1/5)
x = 12^(1/5) + 1
x = 1.64375 + 1
x = 2.64375
:
:
Check:
(2.64375 - 1)^5 =
(1.64375)^5 = 11.9999 ~ 12